OpenStudy (anonymous):

The quadratic equation x2 – 63x + n = 0 has two prime number roots. If you know that the equation has two prime roots, what value of “n” will make the statement true?

OpenStudy (anonymous):

\[ (x-p_1)(x-p_2) = xx-p_1x-p_2x+(-p_1)(-p_2)=x^2-(p_1+p_2)x+p_1p_2 \]

OpenStudy (anonymous):

This means \[ p_1+p_2=63 \]And \[ p_1p_2=n \]

OpenStudy (anonymous):

So, what two prime numbers add up to \(63\)?

OpenStudy (anonymous):

Here are some prime numbers \[ 2,\; 3,\; 5,\; 7,\; 11,\; 13,\; 17,\; 19,\; 23,\; 29,\; 31,\; 37,\; 41,\; 43,\; 47,\; 53,\; 59,\; 61 \]

OpenStudy (anonymous):

Pick two that add up to \(63\).